OpenStudy (anonymous):

A rectangular flower garden if 12 1/2 ft long and 18 2/3 ft wide. What is the perimeter of the garden?

OpenStudy (jdoe0001):

 \(\bf length =l\qquad width=w\\ perimeter = l+w+l+w=2l+2w\implies 2\left(12\frac{1}{2}\right)+2\left(18\frac{2}{3}\right)\)